OpenStudy (itiaax):

Show that (1-i) is the square root of -2i.

jimthompson5910 (jim_thompson5910):

You have to show that \[\Large 1-i = \sqrt{-2i}\] equivalently, you can show that \[\Large (1-i)^2 = -2i\] I squared both sides to get that second equation

OpenStudy (freckles):

Recall (a+b)^2=a^2+2ab+^2 and not a^2+b^2 Just a reminder

OpenStudy (jdoe0001):

@iTiaax any thoughts? or you fainted ?

OpenStudy (itiaax):

@jdoe0001 , It's all good! I had no idea how simple this was :O